OpenStudy (anonymous):

Solve the trigonometric equation for t. 2 sin t-5=-4, for 0≤t≤2π a) π/6, 5π/6 b) π/3, 2π/6 c) π/4, 3π/4 d) π/6, 7π/6 Please & thank you :)

OpenStudy (kainui):

Where are you running into trouble?

OpenStudy (anonymous):

I have no idea where to start. This is an AP Calculus summer assignment but I don't remember this from precalc.

OpenStudy (kainui):

Basically all you're doing here is algebra. Do the opposite of each thing until you get to your variable on one side. The opposite of sine is arcsine.

OpenStudy (anonymous):

2sint = 5-4 2sint = 1 sint = 1/2 For what values of t between 0 to 2pi, we have 1/2.. First value is 30 or pi/6 So, your answer is either A or D.. Now can you tell me what is the other value of t so that sint = 1/2???

OpenStudy (kainui):

I can't tell if that's a sin(t-5) or a sin(t)-5 though.

OpenStudy (anonymous):

Yeah that's what I can't tell either. On the paper it says 2sin t - 5 = -4..

OpenStudy (anonymous):

sin of 5Ï€/6 = 1/2 @waterineyes

OpenStudy (anonymous):

Yes dear you are right sin150 or sin5pi/6 = 1/2...

OpenStudy (anonymous):

So I chose A, correct?

OpenStudy (anonymous):

Thanks to everyone who helped btw! :)

OpenStudy (anonymous):

Welcome buddy...